# Gianluca Monaci          Rafael S. Rezende          Romain Deffayet          Gabriela Csurka Guillaume Bono          Hervé Déjean          Stéphane Clinchant          Christian Wolf

### v1

@@ 5.2 Retrieval allows zero-shot generalization
By retrieving the image goal $\mathbf{g}^{\mathcal{R}}$ from the database, it is possible to apply any ImageNav model to the ObjectNav and Instance-ImageNav tasks. Tab. 6 compares the performance of zero-shot ImageNav architectures in: (i) Gibson ObjectNav, a variant introduced in [1] for ImageNav agents, consisting of $1,000$ episodes over $5$ Gibson scenes and containing the 6 object categories of HM3D ObjectNav, and (ii) HM3D Instance-ImageNav with modified camera intrinsics and robot size to fit the ImageNav configuration. In both cases we use retrieval databases of $50,000$ images, since it is important to retrieve an image goal $\mathbf{g}^{\mathcal{R}}$ as close as possible to the goal, as shown in Sec. 5.4. In Instance-ImageNav we replace the image goal with the closest image in DINOv2 feature space from the retrieval set. In zero-shot ObjectNav we retrieve the top-9 closest images in CLIP space and re-rank them at each step to select as goal $\mathbf{g}^{\mathcal{R}}$ the element with the highest co-visibility with the current observation, and the remaining 8 images constitute the context. This is done using the frozen co-visibility auxiliary loss head used to train DEBiT models. This step is needed since in ObjectNav any instance of a target object is a valid goal.
Retrieval enables the use of DEBiT-b for both tasks, achieving strong performance in line with best existing methods (for reference, DEBiT-b achieves $10$ SR and $3.1$ SPL on the original Instance-ImageNav task). RANa outperforms the feature-matching based method [42] which uses additional depth and odometry sensors and is the current SotA in zero-shot Instance-ImageNav, as well as ZSON (configuration A) [47] in zero-shot Gibson ObjectNav.
@@
ZER [1] 11.3
ZSON A [47] 31.3 12.0 16.1 8.4
Mod-INN [42] Depth, odom 56.1 23.3
@@
Table 6: Zero-shot navigation results using retrieved image goal $\mathbf{g}^{\mathcal{R}}$ for Gibson ObjectNav and HM3D Instance-ImageNav, and comparison to the state of the art.  obtained with camera intrinsics and robot size of the original Instance-ImageNav task specification.
Fig. 4 shows an example of successful zero-shot Instance-ImageNavigation, where our RANa agent first correctly picks from the context the door near the goal ①, and then a view of the room where the target is ②, leading to success. In contrast, the DEBiT-b baseline without context does not see the target inside the room and passes by, exploring the wrong side of the house and failing to complete the episode.
Figure 4: Zero-shot Instance-ImageNavigation episode by RANa and by the DEBiT-b baseline without context (top-right box). The green camera icon indicates the pose of the retrieved image goal $\mathbf{g}^{\mathcal{R}}$, the pink ones indicate the poses of selected context items.
@@ 5.3 Retrieval allows bridging the sim-to-real gap
We evaluate our [Uncaptioned image] RANa-b agent on a small Instance-ImageNav dataset of $200$ episodes where goal images are pictures taken in a real office environment. We manually annotate the positions of $20$ real images in a simulated 3D reconstruction of the environment. We generate a retrieval set $\mathcal{D}$ of $10$k images from uniformly-sampled navigable poses, rendered in simulation using the agent configuration. Simulating navigation from $10$ random starting poses towards each of the $20$ real image goals, RANa reaches $46.5$ success, vs. $40.5$ of DEBiT-b (see inlet Table in Fig. 5). The Figure illustrates a challenging success case (left), where the cactus viewing angle is far from any view the agent could capture, and (right) an interesting failure case, where the retrieval returned as goal $\mathbf{g}^{\mathcal{R}}$ an image of a different instance of the same microwave model at a different position. Interestingly, some retrieved context images depict the correct goal, and RANa appears to be able to exploit this information as it succeeds in $3/10$ episodes featuring this goal.
@@
Figure 5: Navigation in simulation from real goal images, retrieving new goals: 2 episodes shown, 200 episodes evaluated.
5.4 Further ablation studies
All ablations were conducted on RANa-t models based on the DEBiT-t backbone to allow for faster training and evaluation. Additional ablations are reported in Supplementary.
Model Retrieval Encoder Retr. Feature SR SPL
[Uncaptioned image] DEBiT-t 79.9 51.4
[Uncaptioned image] RANa-t DINOv2 Gumbel CroCo Geom. FM 83.0 52.3
[Uncaptioned image] RANa-t DINOv2 Gumbel hwRN18 79.1 48.2
Table 7: [Uncaptioned image]  Impact of Geometric FM on context representation.
Geometric FM provides essential directional information – and plays a crucial role in the effectiveness of RANa. We validate this intuition by training with the same procedure as all other models a RANa-t ImageNav network where the encoders $d$ in Eq. 6 are replaced by a hwRN18 (same as encoder $x$) that takes as $9{\times}112{\times}112$ input the channel-concatenated observation, goal and context, and outputs the context features $\mathbf{e}_{t,n}$, fed to the Gumbel selector. As shown in Tab. 7, this model achieves $79.1$ SR, against $83.0$ SR of RANa-t: in this setting, the agent is not capable of exploiting context information and performance remain at DEBiT-t baseline level, validating the importance of the geometric pre-training done with the CroCo FM in [6].
Retrieval is efficientTab. 8 shows model size, retrieval ($t_{R}$) and inference ($t_{I}$) time for selected DEBiT and RANa models. Retrieval time is low (negligible in non-graph case), could even be further optimized by not retrieving at every step, and scales favorably with the dataset size thanks to efficient approximate nearest neighbor search libraries such as Faiss [22]. Inference time is slightly higher for RANa compared to DEBiT because of encoder $c$ that computes CroCo features for context images, observation and goal. Still, RANa-t is smaller and faster than DEBiT-b while achieving similar performance, suggesting that retrieval augmentation can be a frugal alternative to model scaling.
Context diversity is importantTab. 9 evaluates the behavior of RANa-t ImageNav model when using different strategies to create the context, (i) rand: 8 random elements from the retrieval dataset, (ii) top 8: most similar database images to the goal in DINOv2 feature space, and (iii) top 8 MMR: most similar database images filtered by MMR as in Sec. 4.1. Filtering the context with MMR is important, as the agent relies on diverse context items; removing MMR degrades performance. Other approaches have lower but more constant performance across settings. Similar experiments and behavior is presented for ObjectNav in Supplementary.
Model Retrieval Nr Par $t_{R}(s)$ $t_{I}(s)$ SR SPL
DEBiT-t 27M 0.016 79.3 50.0
DEBiT-b 66M 0.025 83.0 55.6
RANa-t DINO-Gumbel 32M 0.001 0.021 83.0 52.3
RANa-t DINOGraph-Attn 34M 0.055 0.024 84.7 63.7
Table 8: Retrieval and inference runtimes (Tesla-T4-16G GPU).
rand top 8 top 8 MMR
rand 81.9 (51.3) 80.3 (50.4) 81.0 (50.3)
top 8 78.4 (49.0) 81.4 (48.1) 80.7 (51.1)
top 8 MMR 77.0 (47.1) 73.7 (42.4) 83.0 (52.3)
Table 9: [Uncaptioned image]  Impact of retrieval data selection - SR (SPL) for ImageNav with different context building strategies, combinations of train and test conditions.
@@ 6 Conclusion
We propose a general retrieval-augmented navigation agent, trained with RL, that is able to retrieve and act on images stored in a large, global database of robot observations. We leverage pre-trained foundation models to allow the agent to query and process contextual information that can be useful for solving its target task. Augmenting ImageNav and ObjectNav agents with a context retrieval module improves navigation performance in both tasks, despite the goal specifications by different modalities (an image in ImageNav , text in ObjectNav), which showcases the generality of the proposed method. Moreover, we show how to use retrieval across domains and modalities to apply existing agents to new tasks in a zero-shot fashion.
@@
Appendix A Implementation details
A.1 Retrieval dataset creation
@@
Both ImageNav and ObjectNav benchmarks use the Habitat simulator, platform and task definitions [64]. For ImageNav, we use the Gibson dataset [77], consisting of 72 train and 14 eval scenes, and the ImageNav agent configuration: height 1.5m, radius 0.1m, $112{\times}112$ RGB camera with $90^{\circ}$ horizontal field-of-view (hfov) at 1.25m from the ground. For ObjectNav, we use 80 train and 20 eval scenes of the HM3DSem-v0.2 [60, 81] as in [10] and an agent of height 0.88m, radius 0.18m with $112{\times}112$ RGB sensor with $79^{\circ}$ hfov placed at 0.88m from the ground. For each scene we generate a retrieval dataset $\mathcal{D}$ usually containing $1,000$ FPV images as follows: (i) we randomly choose start and goal viewpoints such that the goal is navigable from the start, (ii) we make an agent follow the shortest path from start to goal, and (iii) after each action, we save the agent’s RGB observation as an image in $\mathcal{D}$. We repeat this process until the dataset contains the desired number of images. We randomly choose the distribution of starts and goals such that the agent roughly covers all navigable areas, and that the path between pairs of viewpoints are not too close to one another to simulate a prior offline exploration of the scene.
A.2 RANa agent architecture
@@
In this section, we further explain some of the architecture choices made for our method. Observation encoderx processes the RGB input image $\mathbf{x}_{t}\in\mathbb{R}^{3\times 112\times 112}$. It is implemented with a half-width ResNet-18 [31] and generates an output feature $\tilde{\mathbf{x}}_{t}\in\mathbb{R}^{512}$.
@@
Goal comparison functiong compares the observation $\mathbf{x}_{t}$ to the goal $\mathbf{g}$. For ImageNav we use the frozen binocular encoder of DEBiT [6], based on the Geometric Foundation Model (FM) CroCo [74]. This encoder is a Siamese encoder $E$ applied to both $\mathbf{x}_{t}$ and $\mathbf{g}$, and a decoder $D$ which combines the output of the two encoders. Both $E$ and $D$ are implemented with a ViT architecture with self-attention layers, and $D$ also adds cross-attention. The output of the decoder $D$ is further compressed by a fully connected layer (frozen, from the DEBiT model) that projects the flattened output of $D$ into $\tilde{\mathbf{g}}_{t}\in\mathbb{R}^{3136}$. For details about the binocular encoder $g$ we refer to [6] and [74].
@@
In ObjectNav we use OpenCLIP [35] to encode the visual observation $\mathbf{x}_{t}$ and the textual goal $\mathbf{g}$. This process produces two features $cl_{V}(\mathbf{x}_{t})\in\mathbb{R}^{512}$ and $cl_{T}(\mathbf{g})\in\mathbb{R}^{512}$, which are concatenated to generate the output $\tilde{\mathbf{g}}_{t}\in\mathbb{R}^{1024}$.
@@
Action embeddingl encodes the previous action $\mathbf{a}_{t-1}$ into a $32$-dimensional feature vector $\tilde{\mathbf{a}}_{t}$ and it is taken from the DEBiT [6] model. For ImageNav, we load DEBiT action encoder and freeze it during training, while for ObjectNav, we trained it from scratch, since the two tasks do not share a common action space. Indeed, the task definition of ObjectNav requires two additional actions, LOOK UP and LOOK DOWN.
@@
Gumbel context encoderVariant 1 of the trainable context encoder c (Sec. 3.1 ) selects one element $\tilde{\mathbf{c}}_{t}$ among the context features $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ based on a learned relevance estimate $\{\alpha_{n}\}_{n=1}^{N}$. The retrieved context images are compared to the goal image (if available) and to the observation by the same frozen binocular encoder of DEBiT described above, leading to the $N$ context features $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ each of dimension $\mathbb{R}^{D}$, where $D=6272$ for ImageNav or Instance-ImageNav (when goal image is available), and $D=3136$ for ObjectNav. A fully connected layer $\texttt{Linear}(D,1)$ (using a syntax inspired by PyTorch) estimates the relevance $\{\alpha_{n}\}_{n=1}^{N}$ for each context element and the item with largest relevance is selected using the Gumbel soft-max sampler [37]. In the ImageNav or Instance-ImageNav cases, only the binocular feature representing the comparison between observation and context item is selected, so in all cases the context feature $\tilde{\mathbf{c}}_{t}\in\mathbb{R}^{3136}$.
Attention-based context encoderVariant 2 of the trainable context encoder creates a context feature by computing the self-attention over the context elements $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ computed as described in the previous paragraph. In this case the context encoder c first reduces the dimensionality of each context feature to $\mathbb{R}^{128}$ through a fully connected layer $\texttt{Linear}(D,128)$ ($D=6272$ for ImageNav or Instance-ImageNav, and $D=3136$ for ObjectNav), then processes them with a standard two-layer transformer encoder network [72] with $4$ attention heads and feedforward dimension $256$. The $N$ context features are concatenated into a feature vector of size $\mathbb{R}^{N\times 128}$ and processed by a 2-layer $\texttt{MLP}(128,3136)$ with hidden dimension $128$ and output $3136$, such that the context encoder output $\tilde{\mathbf{c}}_{t}$ has the same dimension of that of the Gumbel context encoder.
@@
Recurrent memory – The flatten vectors $\tilde{\mathbf{x}}_{t}$, $\tilde{\mathbf{g}}_{t}$, $\tilde{\mathbf{a}}_{t}$ and $\tilde{\mathbf{c}}_{t}$ are concatenated and fed to a single-layer GRU [18] with hidden state $\mathbf{h}_{t}\in\mathbb{R}^{512}$. This GRU is an “extended” version of the baseline agent’s GRU without context, which is finetuned. However, since RANa concatenates the context feature $\tilde{\mathbf{c}}_{t}$ to the other inputs, we cannot directly finetune the GRU as is. We modify the GRU of the base model by padding the input weight matrices of the first layer, $W^{0}_{ir},W^{0}_{iz},W^{0}_{in}$ (and corresponding bias vectors $b^{0}_{ir},b^{0}_{iz},b^{0}_{in}$) along the input dimension with random values. The remaining GRU matrices are initialized to the values of the baseline checkpoints and finetuned from there.
@@
A.3 DINOv2-Graph
In this section we provide details about the construction of the DINOv2-Graph context. The graph is built from the affinity matrix $\Omega_{\mathcal{D}}$ containing the similarity between DINOv2 [55] features of all pairs of images in $\mathcal{D}$, where the nodes are images $\mathbf{x}_{i}^{\mathcal{D}}\in\mathcal{D}$ and edges are weighted by similarity. We use Dijkstra’s algorithm [20] (from the SciPy python package) to select the shortest path between two nodes.

### v2

@@ 5.2 Retrieval allows zero-shot generalization
ZSEL (Al-Halah et al., 2022) 11.3
ZSON A (Majumdar et al., 2022) 31.3 12.0 16.1 8.4
Mod-INN (Krantz et al., 2023) Depth, odom 56.1 23.3
@@
Table 8: Zero-shot navigation results using retrieved image goal $\mathbf{g}^{\mathcal{R}}$ for Gibson ObjectNav and HM3D Instance-ImageNav, and comparison to state-of-the-art.  obtained with camera intrinsics and robot size of the original Instance-ImageNav task specification.
By retrieving the image goal $\mathbf{g}^{\mathcal{R}}$ from the database, it is possible to apply any ImageNav model to the ObjectNav and Instance-ImageNav tasks. Table˜8 compares the performance of zero-shot architectures in: (i) Gibson ObjectNav, a variant introduced in Al-Halah et al. (2022) for ImageNav agents, consisting of 1,0001,0001 , 000 episodes over 555 Gibson scenes and containing the 6 object categories of HM3D ObjectNav (chair, bed, plant, sofa, tv, toilet), and (ii) HM3D Instance-ImageNav (Krantz et al., 2023) with modified camera intrinsics and robot size to fit the ImageNav configuration. In both cases we use retrieval databases of 50,00050,00050 , 000 images, as it is important to retrieve an image goal $\mathbf{g}^{\mathcal{R}}$ as close as possible to the real goal to avoid stopping too far from it. We ablate this parameter in Appendix˜D and show that RANa already achieves SOTA zero-shot Instance-ImageNav performance with as little as 5,0005,0005 , 000 retrieval images. In Instance-ImageNav we replace the image goal with the closest image in DINOv2 feature space from the retrieval set. In zero-shot ObjectNav, where the goal is defined as an object category, we retrieve the top-9 closest dataset images in CLIP space (Ilharco et al., 2021) and re-rank them at each step to select as goal $\mathbf{g}^{\mathcal{R}}$ the element with the highest co-visibility with the current observation, and the remaining 8 images constitute the context. This is done using the frozen co-visibility auxiliary loss head used to train DEBiT models. This step is needed since in ObjectNav any instance of a target object is a valid goal.
Retrieval enables the use of DEBiT-b for both tasks, achieving strong performance in line with best existing methods (for reference, zero-shot DEBiT-b without retrieved goal image achieves 101010 SR and 3.13.13.1 SPL on Instance-ImageNav). RANa outperforms the feature-matching based method of Krantz et al. (2023) which uses additional depth and odometry sensors and is the current SOTA in zero-shot Instance-ImageNav, as well as ZSON (configuration A) of Majumdar et al. (2022) in zero-shot Gibson ObjectNav.
Fig.˜5 shows an example of successful zero-shot Instance-ImageNavigation, where our RANa agent first correctly picks from the context the door near the goal ①, and then a view of the room where the target is ②, leading to success. In contrast, the DEBiT-b baseline without context does not see the target inside the room and passes by, exploring the wrong side of the house and failing to complete the episode.
Figure 5: Zero-shot Instance-ImageNav episode by RANa-b and by DEBiT-b baseline (top-right box). The green camera indicates the pose of the retrieved image goal $\mathbf{g}^{\mathcal{R}}$, the pink ones indicate the poses of selected context items.
@@ 5.3 Retrieval allows bridging the sim-to-real gap
We evaluate our [Uncaptioned image] RANa-b agent on a small Instance-ImageNav dataset of 200200200 episodes where goal images are taken in a real office environment. We manually annotate the positions of 202020 real images in a simulated 3D reconstruction of the environment. We generate a retrieval set $\mathcal{D}$ of 101010k images from uniformly-sampled navigable poses, rendered in simulation using the agent configuration. Simulating navigation from 101010 random starting poses towards each of the 202020 real image goals, RANa reaches 46.546.546.5 success, vs. 40.540.540.5 of DEBiT-b (see inlet Table in Fig.˜6). The Figure illustrates a challenging success case (left), where the cactus viewing angle is far from any view the agent could capture, and (right) an interesting failure case, where the retrieval returned as goal $\mathbf{g}^{\mathcal{R}}$ an image of a different instance of the same microwave model at a different position. Interestingly, some retrieved context images depict the correct goal, and RANa appears to be able to exploit this information as it succeeds in 3/103/103 / 10 episodes featuring this goal.
@@
Figure 6: Navigation with real goal images, retrieving new goals: 2 episodes shown, 200 episodes evaluated.
6 Extension to ObjectNav
We extend the proposed approach to ObjectNav, where the goal is a textual label of an object category, by slightly modifying the RANa architecture. Since ObjectNav requires semantic understanding of the scene and common-sense reasoning, a large body of research builds semantic maps (Gadre et al., 2022; Chaplot et al., 2020a) and/or leverages an LLM to guide navigation based on a scene’s common sense semantics (Yu et al., 2023; Cai et al., 2024; Kuang et al., 2024; Yin et al., 2024). Few approaches address this task in a zero-shot fashion, training their agents without ObjectNav rewards and leveraging CLIP encoders for goal representation (Sun et al., 2024b; Majumdar et al., 2022; Gadre et al., 2023). Here we do not aim to achieve state-of-the-art results, as it is unlikely for our simple approach to compete with methods using LLMs and VLMs with billion of parameters. Instead we want to showcase the effectiveness and flexibility of our method.
In order to do that, we simply modify few components of the RANa-t static architecture: to account for the textual nature of the goal we use the multi-modal FM OpenCLIP (Ilharco et al., 2021) to compute the goal representation $\tilde{\mathbf{g}}_{t}$ and the similarity measure used for retrieval. Besides, ObjectNav has two additional actions, LOOK UP and LOOK DOWN, that are added to the action space. We train and test this agent, as well as a baseline not augmented with retrieval (we refer to it as DEBiT-t ObjectNav), on the HM3DSem dataset (Cai et al., 2024; Ramakrishnan et al., 2021) with 6 object categories (chair, bed, plant, sofa, tv, toilet) using the standard Habitat ObjectNav task definition and following the same procedure described earlier.
Table˜9 compares the results of RANa-t and DEBiT-t ObjectNav models with competing approaches. RANa can exploit retrieved context also in this setting and improves the DEBiT baseline SR by almost +10+10+ 10 points. Compared to methods with similar sensor settings (only one RGB camera), RANa-t outperforms previous SOTA by +9.7+9.7+ 9.7 on SR and +7.7+7.7+ 7.7 on SPL. We also compare RANa-t to methods that use extra sensors, such as depth camera, GPS+compass, or 6 camera sensors, as well as LLM-guided navigation. Here RANa-t shows competitive results, outperforming several strong baselines such as ESC (Zhou et al., 2023) and PixNav (Cai et al., 2024).
Model Extra sensors w/ map w/ LLM SR SPL
PixNav (Cai et al., 2024) Pano RGB 37.9 20.5
ESC (Zhou et al., 2023) Depth, odom 39.2 22.3
L3MVN (Yu et al., 2023) Depth, odom 54.2 25.5
CoW (Gadre et al., 2022) Depth, odom 6.1 3.9
GoW (Gadre et al., 2023) Depth, odom 32.0 18.1
ProcTHOR (Deitke et al., 2022) Depth, odom 13.2 7.7
OVRL-v2 (Yadav et al., 2023a) Odom 64.7 28.1
ZSON (Majumdar et al., 2022) 25.5 12.6
PSL (Sun et al., 2024b) 42.4 19.2
DEBiT-t ObjectNav 42.9 23.0
RANa-t ObjectNav 52.1 26.9
Table 9: DEBiT-t and RANa-t ObjectNav compared to state-of-the-art methods.
@@ 7 Conclusion
We propose a general retrieval-augmented navigation agent, trained with RL, that is able to retrieve and act on images stored in a large, global database of robot observations. We leverage pre-trained foundation models to allow the agent to query and process contextual information that can be useful for solving its target task. Augmenting a SOTA ImageNav agent with a context retrieval module improves navigation performance while adding minimal overhead, which showcases the flexibility and effectiveness of the proposed method. Moreover, we demonstrate the approach for an ObjectNav agent, and we show how to use retrieval across domains and modalities to apply existing agents to new tasks in a zero-shot fashion.
@@
Appendix A Retrieval dataset creation
We use the Habitat simulator, platform and task definitions from Savva et al. (2019). For ImageNav, we use the Gibson dataset (Xia et al., 2018), consisting of 72 train and 14 eval scenes, and the ImageNav agent configuration: height 1.5m, radius 0.1m, ${\times}112$ 112 RGB camera with $90^{\circ}$ horizontal field-of-view (hfov) at 1.25m from the ground. To train and test the ObjectNav variant in Section˜6, we use 80 train and 20 eval scenes of the HM3DSem-v0.2 (Ramakrishnan et al., 2021; Yadav et al., 2023c) as in Cai et al. (2024) and an agent of height 0.88m, radius 0.18m with ${\times}112$ 112 RGB sensor with $79^{\circ}$ hfov placed at 0.88m from the ground.
For each scene, we generate a retrieval dataset $\mathcal{D}$ usually containing 1,0001,0001 , 000 FPV images as follows:
  1. (i)
    we randomly choose start and goal viewpoints such that the goal is navigable from the start,
  2. (ii)
    we make an agent follow the shortest path from start to goal, and
  3. (iii)
    after each action, we save the agent’s RGB observation as an image in $\mathcal{D}$.
@@ Appendix A Retrieval dataset creation
We repeat this process until the dataset contains the desired number of images. We randomly choose the distribution of starts and goals such that the agent roughly covers all navigable areas, and that the path between pairs of viewpoints are not too close to one another to simulate a prior offline exploration of the scene.
Appendix B RANa agent architecture
In this section we detail some of the architecture choices made for our method.
@@ Appendix B RANa agent architecture
Observation encoderx processes the RGB input image $\mathbf{x}_{t}\in\mathbb{R}^{3\times 112\times 112}$. It is implemented with a half-width ResNet-18 (He et al., 2016) and generates an output feature $\tilde{\mathbf{x}}_{t}\in\mathbb{R}^{512}$.
@@
Goal comparison functiong compares the observation $\mathbf{x}_{t}$ to the goal $\mathbf{g}$. For ImageNav we use the frozen binocular encoder of DEBiT (Bono et al., 2024a), based on the Geometric Foundation Model (FM) CroCo (Weinzaepfel et al., 2022). This encoder is a Siamese encoder $E$ applied to both $\mathbf{x}_{t}$ and $\mathbf{g}$, and a decoder $D$ which combines the output of the two encoders. Both $E$ and $D$ are implemented with a ViT architecture with self-attention layers, and $D$ also adds cross-attention. The output of the decoder $D$ is further compressed by a fully connected layer (frozen, from the DEBiT model) that projects the flattened output of $D$ into $\tilde{\mathbf{g}}_{t}\in\mathbb{R}^{3136}$. For details about the binocular encoder $g$ we refer to Bono et al. (2024a) and Weinzaepfel et al. (2022).
@@
In ObjectNav we use OpenCLIP (Ilharco et al., 2021) to encode the visual observation $\mathbf{x}_{t}$ and the textual goal $\mathbf{g}$. This process produces two features $cl_{V}(\mathbf{x}_{t})\in\mathbb{R}^{512}$ and $cl_{T}(\mathbf{g})\in\mathbb{R}^{512}$, which are concatenated to generate the output $\tilde{\mathbf{g}}_{t}\in\mathbb{R}^{1024}$.
@@
Action embeddingl encodes the previous action $\mathbf{a}_{t-1}$ into a 323232-dimensional feature vector $\tilde{\mathbf{a}}_{t}$ and it is taken from the DEBiT (Bono et al., 2024a) model. For ImageNav, we load DEBiT action encoder and freeze it during training, while for ObjectNav, we trained it from scratch, since the two tasks do not share a common action space. Indeed, the task definition of ObjectNav requires two additional actions, LOOK UP and LOOK DOWN.
@@
Gumbel context encoderVariant 1 of the trainable context encoder c (Section˜3.1) selects one element $\tilde{\mathbf{c}}_{t}$ among the context features $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ based on a learned relevance estimate $\{\alpha_{n}\}_{n=1}^{N}$. The retrieved context images are compared to the goal image (if available) and to the observation by the same frozen binocular encoder of DEBiT described above, leading to the $N$ context features $\{\mathbf{e}_{t,n}\}_{n=1}^{N}$ each of dimension $\mathbb{R}^{D}$, where $D=6272$ for ImageNav or Instance-ImageNav (when goal image is available), and $D=3136$ for ObjectNav. A fully connected layer $\texttt{Linear}(D,1)$ (using a syntax inspired by PyTorch) estimates the relevance $\{\alpha_{n}\}_{n=1}^{N}$ for each context element and the item with the largest relevance is selected using the Gumbel soft-max sampler (Jang et al., 2017). In the ImageNav or Instance-ImageNav cases, only the binocular feature representing the comparison between observation and context items is selected, so in all cases the context feature $\tilde{\mathbf{c}}_{t}\in\mathbb{R}^{3136}$.
[…]
@@
Recurrent memory – The flatten vectors $\tilde{\mathbf{x}}_{t}$, $\tilde{\mathbf{g}}_{t}$, $\tilde{\mathbf{a}}_{t}$ and $\tilde{\mathbf{c}}_{t}$ are concatenated and fed to a single-layer GRU (Cho et al., 2014) with hidden state $\mathbf{h}_{t}\in\mathbb{R}^{512}$. This GRU is an “extended” version of the baseline agent’s GRU without context, which is finetuned. However, since RANa concatenates the context feature $\tilde{\mathbf{c}}_{t}$ to the other inputs, we cannot directly finetune the GRU as is. We modify the GRU of the base model by padding the input weight matrices of the first layer, $W^{0}_{ir},W^{0}_{iz},W^{0}_{in}$ (and corresponding bias vectors $b^{0}_{ir},b^{0}_{iz},b^{0}_{in}$) along the input dimension with random values. The remaining GRU matrices are initialized to the values of the baseline checkpoints and finetuned from there.
@@
Appendix C DINOv2-Graph
In this section we provide details about the construction of the DINOv2-Graph context. The graph is built from the affinity matrix $\Omega_{\mathcal{D}}$ containing the similarity between DINOv2 (Oquab et al., 2024) features of all pairs of images in $\mathcal{D}$, where the nodes are images $\mathbf{x}_{i}^{\mathcal{D}}\in\mathcal{D}$ and edges are weighted by similarity. We use Dijkstra’s algorithm (Dijkstra, 1956) (from the SciPy python package) to select the shortest path between two nodes.